# Stochastic Semi-Gradient Descent for Learning Mean Field Games with Population-Aware Function Approximation

### v1

@@ Abstract
Mean field games (MFGs) model the interactions within a large-population multi-agent system using the population distribution. Traditional learning methods for MFGs are based on fixed-point iteration (FPI), which calculates best responses and induced population distribution separately and sequentially. However, FPI-type methods suffer from inefficiency and instability, due to oscillations caused by the forward-backward procedure. This paper considers an online learning method for MFGs, where an agent updates its policy and population estimates simultaneously and fully asynchronously, resulting in a simple stochastic gradient descent (SGD) type method called SemiSGD. Not only does SemiSGD exhibit numerical stability and efficiency, but it also provides a novel perspective by treating the value function and population distribution as a unified parameter. We theoretically show that SemiSGD directs this unified parameter along a descent direction to the mean field equilibrium. Motivated by this perspective, we develop a linear function approximation (LFA) for both the value function and the population distribution, resulting in the first population-aware LFA for MFGs on continuous state-action space. Finite-time convergence and approximation error analysis are provided for SemiSGD equipped with population-aware LFA.
@@ 1 Introduction
Mean field games (MFGs) [17, 21] offer a tractable framework for modeling multi-agent systems with a large homogeneous population. In MFGs, the impact of other agents on a particular agent is encapsulated by a population mass (or mean field), providing a reliable approximation of actual interactions among agents when the number of agents is large. Consequently, understanding the population is fundamental in MFGs, as learning these games entails considering both the agent’s policy and the population dynamics.
@@
Prior work on learning MFGs has mainly focused on fixed-point iteration (FPI) methods and their variations, which sequentially calculate the best response w.r.t. a fixed population, and the induced population distribution (or measure) w.r.t. a fixed policy [15, 14, 12, 30, 24]. However, FPI-type methods face several limitations: 1) Calculating best responses (backward) and induced population distributions (forward) typically involves distinct iterative processes. These processes are implemented separately and executed sequentially, hindering parallel computing and potentially increasing the overall computational burden. 2) FPI methods suffer from instability. Fixing a population measure far from the equilibrium measure can lead to a corresponding best response that is also far from the equilibrium policy, resulting in slow convergence. Moreover, drastic differences between best responses and induced population measures in consecutive iterations can cause oscillations in the learning process, a phenomenon commonly observed in practice (see, e.g., Figure 1(a) and [12]).
(a) Convergence performance.
[…]
(b) Learning dynamics.
Figure 1: FPI versus SGD
@@
This work delves into the rapidly growing field of online learning for MFGs, where an online agent interacts with the environment and gathers observations on a Markov chain to update its policy and population estimate. Notably, this Markov chain is parameterized by the policy and a population distribution, which collectively determine the transition kernel. Thus, the goal of finding a mean field equilibrium (MFE) can be translated into finding an optimal parameter, which corresponds to the fixed point of the Bellman operator and the transition operator. This perspective unlocks a large toolbox of stochastic parameter approximation methods for learning MFGs, such as simple stochastic gradient descent (SGD)-type methods. Specifically, we treat the policy and the population as a unified parameter and update their estimates fully asynchronously using the same samples, thereby eliminating the forward-backward process.
We term our method SemiSGD, as it uses semi-gradients in lieu of actual gradients. Albeit simple, an SGD-type procedure raises several questions: 1) Simultaneously and asynchronously updating the behavior policy and the population estimate creates a strong coupling between the two, rendering the underlying Markov chain non-stationary. Online FPI avoids this issue by fixing one estimate while updating the other. How can we analyze this coupling and the non-stationary process to ensure the convergence of SemiSGD? 2) How does SemiSGD, following a non-stationary Markov chain, converge? Futhermore, how can it address the instability and oscillation issues of FPI-type methods?
For the first question, note that the Bellman and transition operators are contractive. Thus, with sufficiently smooth dynamics, the non-stationarity is dominated by the contraction property. Remarkably, we discover that SemiSGD indeed follows a descent direction towards the MFE (Lemma 1), provided that the dynamics are Lipschitz continuous with samll Lipschitz constants. By following a descent direction, updating parameters incrementally, and dropping the forward-backward procedure, SemiSGD achieves automatic stabilization without additional mechanics (see Figure 1(a) for a numerical example and Figure 1(b) for an intuitive illustration).
Treating the policy and population as a unified parameter also illuminates the gap in learning MFGs on continuous state-action spaces: linear function approximation (LFA), a technique widely employed in learning value functions on continuous state-action spaces [26, 19], can be applied to the population measure as well. As a framework for infinite-player games, common MFGs are often on infinite or continuous state-action spaces, such as autonomous driving [9], flocking [31] and crowd modelling [20, 7]. Equipped with population-aware LFA (PA-LFA), SemiSGD is capable of learning MFGs on continuous state-action spaces, fully online and model-free with strong theoretical guarantees.
@@
Main results. We highlight some contributions and novelties of this work.
@@
  • We propose SemiSGD, a simple online SGD-type method for learning MFGs. We innovatively treat the policy and the population as a unified parameter. Algorithmically, we update both simultaneously and asynchronously using the same online observations and learning rate, thus eliminating the forward-backward process typical of FPI methods. SemiSGD enjoys automatic stabilization without additional mechanisms such as fictitious play or online mirror descent (Figure 1(a)). Theoretically, we show that the unified parameter in SemiSGD follows a descent direction towards the MFE, whereas neither the policy nor population alone is guaranteed to do so (Lemma 1).
@@
  • Sample complexity and approximation error analysis are provided for SemiSGD with PA-LFA. We novelly regard the learning process as a non-stationary Markov decision process parameterized by the unified parameter. This perspective enables a straightforward SGD-type analysis, elegantly handling the strong coupling between the policy and population and offering insights into the learning dynamics of SemiSGD.
  • We conduct various experiments to demonstrate the properties of our methods: 1) For continuous state-action spaces, PA-LFA is more efficient and accurate than discretization; 2) SemiSGD is inherently smoother and automatically stabilizes without additional mechanics; 3) By eliminating the forward-backward procedure, SemiSGD is significantly faster than other FPI-type methods.
Table 1: Comparison of learning methods for MFGs
Reference Algo. type Additional mechanism Oracle- free Function approx. Sample complex.
Population Policy
[14, 15, 35] FPI - - - - -
[30, 24] FPI FP & OMD - - - -
[1, 12] FPI regularization - - - -
[4, 2] FPI multi-timescale - - -
[37] FPI FP - - $\widetilde{O}(\epsilon^{-2})$
[36] FPI two-timescale - - $O(\epsilon^{-4})$
[25] FPI two-timescale - - $\widetilde{O}(\epsilon^{-5})$
[16] FPI - - - $\widetilde{O}(\epsilon^{-2})$
This paper SGD - $\widetilde{O}(\epsilon^{-2})$
Related work. There is a growing body of literature on learning MFGs (summarized in Table. 1). Readers can refer to [23, 13] for a comprehensive review. To address the instability issue of FPI, researchers have proposed various methods to stabilize the learning process, broadly categorized into two classes: 1) regularization [1, 12]; 2) fictitious play (FP) [30, 8, 35]; and 3) online mirror descent (OMD) [29, 24]. A multi-timescale [2, 4, 36, 25] update rule is utilized for the population and value function in the forward-backward process. By fixing one element (or make its update much slower) while updating the other, these methods circumvent the strong coupling between the value function and population measure, and thus inherit the convergence properties of FPI-type methods. To deal with large state and action spaces and obtain strong theoretical guarantees, [25, 16] consider linear function approximation (LFA) [6, 26, 19] for the policy learning in the backward process of MFGs. The LFA scheme is not incorporated into the coupled forward-backward process, which renders the learning algorithm still a FPI-type method, and sample complexity results fail to capture the population estimate due to the requirement of population oracles in the forward process.
@@ 2.1 Revisit online learning for MFGs on finite state-action spaces
We consider an infinite-horizon Markov decision process (MDP) denoted by $(\mathcal{S},\mathcal{A},r,P,\gamma)$, with the state space $\mathcal{S}$ and action space $\mathcal{A}$ being finite. In MFGs, the reward function $r$ and transition kernel $P$ depend on the population distribution over the state space. Specifically, for a given state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ and population distribution $\mu\in\mathcal{D}(\mathcal{S})$, $r(s,a,\mu)$ and $P(s^{\prime}\,|\,s,a,\mu)$ denote the reward received and the probability of transitioning to state $s^{\prime}\in\mathcal{S}$. $\gamma$ is the discount factor.
An agent in an MFG aims to find a policy $\pi$, which maps a state to a distribution on the action space determining the agent’s actions, that maximizes its expected cumulative discounted reward while interacting with the population. We utilize a value-based approach to calculate policies and assume access to a policy operator $\Gamma_{\pi}$ that returns a policy based on an (action-)value function $Q\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. We write $\pi_{Q}\coloneqq\Gamma_{\pi}(Q)$. To define the learning goal, the mean field equilibrium (MFE), we first introduce the following two operators for any value functions $Q,Q^{\prime}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ and population measures $M,M^{\prime}\in\mathcal{D}(\mathcal{S})$:
@@
A typical FPI-type method (approximately) calculates the fixed point of the Bellman operator given the current population distribution, i.e., the best response, and the fixed point of the transition operator given the current value function, i.e., the induced population distribution, alternately. Under certain conditions, FPI converges to the MFE [15].
This work focuses on online learning (model-free and stochastic) methods for MFGs. In such methods, an agent maintains estimates about the value function and population measure, and uses online observations to update its estimates. For the value function, temporal difference (TD) methods are widely used [32]. Given an online observation tuple $O=(s,a,s^{\prime},a^{\prime})$ and step-size $\alpha$, on-policy TD (or SARSA [32]), updates the Q-value function as follows:
@@
For the population distribution, Monte Carlo (MC) sampling [22] is a common choice. Given a new state sample $s^{\prime}$ and using a general step-size $\alpha$, the population measure is updated as follows:
@@
TD and MC updates are widely used in online learning methods for MFGs [2, 36, 37]. Fixing the policy and population measure (i.e., fixing the transition kernel), TD converges to the optimal value function, and MC converges to the induced population distribution.
@@ 2.2 Stochastic semi-gradient descent
Prior work on online learning for MFGs either uses two-layered loops with the policy and/or population measure fixed in the inner loop [34, 37], or uses a two-timescale to update the policy and population measure in different rates [2, 3], essentially following a FPI-type forward-backward process. In contrast, we propose simply combining Equations 2 and 3 in one pass, resulting in a simple SGD-type method that converges to the MFE efficiently.
@@
To elaborate, $\mathfrak{g}$ in Equation 2 is referred as semi-gradients, as it is not a real gradient of any loss function, but provides a similar descent direction to the stationary point, i.e., the zero point of $\mathfrak{g}$ [^1] With a slight abuse of notation, we use a single operator $\mathfrak{g}$ to return semi-gradients throughout the paper, for both the action-value function and population measure with or without LFA; it should be clear from the context and its arguments which one it refers to. Please refer to Section 4 for the derivation of the population measure semi-gradient in Equation 3. Consequently, we propose updating the value function and population measure estimates simultaneously and asynchronously using semi-gradients in one pass. We term our method SemiSGD and present it in Algorithm 1. In Algorithm 1, we denote $(Q_{t};M_{t})$ as the concatenated representation of the value function and population measure, as they are both vectors for finite state-action spaces. Similarly, $\mathfrak{g}$ represents the concatenation of the semi-gradients for the value function and population measure.
@@
Algorithm 1 is an online single-loop method that completely removes the need for a forward-backward process. Not only it updates the value function and population measure estimates in an SGD-like manner, the following lemma shows that the a descent direction is provided by the mean-path semi-gradient, defined as $\bar{\mathfrak{g}}_{(Q,M)}(\cdot)\coloneqq\mathbb{E}_{(Q,M)}\mathfrak{g}(\cdot;O)$, where $O$ is the online observation tuple following the steady distribution induced by the policy and transition kernel determined by $Q$ and $M$.
Lemma 1 (Descent direction).
Suppose $(Q_{*};M_{*})$ is an MFE and Assumptions 1, 2 and 3 hold. [^2] Section 3 demonstrates linear MFGs encompass finite MFGs (Example 1). Thus, our analysis focuses on linear MFGs, with results subsequently reduced to finite MFGs in Appendix H. As a result, in this section we sometimes refer to the analysis for linear MFGs in later sections; in such cases, finite MFGs are considered as linear MFGs. For any value function $Q$ and population measure $M$, with $\Delta Q\coloneqq Q-Q_{*}$ and $\Delta M\coloneqq M-M_{*}$, we have
$\Delta M,\bar{\mathfrak{g}}_{{(Q,M)}}(M)\right>\leq-\|\Delta M\|^{2}+\sigma L\|\Delta M\|(\|\Delta Q\|+\|\Delta M\|),$ (4)
$\Delta Q,\bar{\mathfrak{g}}_{{(Q,M)}}(Q)\right>\leq-(1-\gamma)\lambda\|\Delta Q\|^{2}+(\sigma LH+L_{r})\|\Delta Q\|(\|\Delta Q\|+\|\Delta M\|),$
where $\lambda\coloneqq\min_{s,a}M_{*}(s)\pi_{Q_{*}}(a\,|\,s)$, $H=O(\sqrt{|\mathcal{S}||\mathcal{A}|}R/(1-\gamma))$, and other constants are defined in Assumptions 1, 2 and 3. Neither $-\bar{\mathfrak{g}}_{{(Q,M)}}(M)$ nor $-\bar{\mathfrak{g}}_{{(Q,M)}}(Q)$ is guaranteed to be a descent direction, no matter how small the Lipschitz constants are. However, if $3\sigma LH+L_{r}\leq(1-\gamma)\lambda$, it holds that
${(Q;M)}-{(Q_{*};M_{*})},\bar{\mathfrak{g}}_{{(Q,M)}}({(Q;M)})\right>\leq-\frac{1}{2}(1-\gamma)\lambda\|{(Q;M)}-{(Q_{*};M_{*})}\|^{2}.$ (5)
@@
A generalized version of Lemma 1 is proved in Section F.1 for linear MFGs. We now discuss some key insights from Lemma 1.
Descent direction. The concatenated semi-gradient points a descent direction for the concatenated estimate $(Q;M)$, while neither the value function nor population measure alone is guaranteed to follow a descent direction. See Figure 9 for an illustrative example. Remarkably, we will show that Algorithm 1 follows the direction of $-\bar{\mathfrak{g}}$ with higher-order noises (Theorem 1). Therefore, Lemma 1 argues that SemiSGD generally follows a descent direction to the MFE.
Explaining oscillations in FPI. Lemma 1 also gives an intuitive explanation for the instability in FPI-type methods. When calculating the best response, if FPI fixes a population measure far from the equilibrium measure (large $\|M-M_{*}\|$), the semi-gradient of the value function may not give a descent direction, potentially driving the best response away from the optimal policy. Similar oscillations can occur when FPI calculates the induced population distribution.
Incremental update by design. To stabilize FPI, fictitious play (FP) and online mirror descent (OMD) incrementally update the policy or population measure [24], rather than switching to entirely new estimates. As an SGD-type method, SemiSGD naturally updates the estimates incrementally, providing another layer of stabilization, and thus does not require those additional mechanisms.
Learn MFG on a non-stationary Markov chain. SemiSGD samples a non-stationary Markovian trajectory, as its transition kernel depends on the changing value function and population measure. Generally, rapidly changing Markov chains exhibit chattering [39]. Thus, Lemma 1 requires that the MDP components be Lipschitz continuous with sufficiently small Lipschitz constants. The proof (Section F.1) shows that this descent direction arises from the contraction properties of the Bellman and transition operators. When the MDP is smooth enough, non-stationarity is dominated by this contraction. The contraction property of the operators translates to the attraction property of their fixed point, i.e., the MFE, which corresponds to a stationary parameter where the mean-field semi-gradient is zero (Proposition 2). This fact again underscores that the landscape of an MFG is best described by the unified parameter, making it more natural to learn this parameter directly.

### v2

@@ Abstract
Mean field games (MFGs) model interactions in large-population multi-agent systems through population distributions. Traditional learning methods for MFGs are based on fixed-point iteration (FPI), where policy updates and induced population distributions are computed separately and sequentially. However, FPI-type methods may suffer from inefficiency and instability due to potential oscillations caused by this forward-backward procedure. In this work, we propose a novel perspective that treats the policy and population as a unified parameter controlling the game dynamics. By applying stochastic parameter approximation to this unified parameter, we develop SemiSGD, a simple stochastic gradient descent (SGD)-type method, where an agent updates its policy and population estimates simultaneously and fully asynchronously. Building on this perspective, we further apply linear function approximation (LFA) to the unified parameter, resulting in the first population-aware LFA (PA-LFA) for learning MFGs on continuous state-action spaces. A comprehensive finite-time convergence analysis is provided for SemiSGD with PA-LFA, including its convergence to the equilibrium for linear MFGs—a class of MFGs with a linear structure concerning the population—under the standard contractivity condition, and to a neighborhood of the equilibrium under a more practical condition. We also characterize the approximation error for non-linear MFGs. We validate our theoretical findings with six experiments on three MFGs.
@@ 1 Introduction
Mean field games (MFGs) (Huang et al., 2006; Lasry & Lions, 2007) offer a tractable framework for modeling multi-agent systems with a large homogeneous population. In MFGs, the impact of other agents on a particular agent is encapsulated by a population mass (or mean field), providing a reliable approximation of actual interactions among agents when the number of agents is large. Consequently, understanding the population is fundamental in MFGs, as learning these games entails considering both the agent’s policy and the population dynamics.
@@
Prior work on learning MFGs has mainly focused on fixed-point iteration (FPI) methods and their variations, which is characterized by a forward-backward procedure that alternately calculate the policy update w.r.t. a fixed population and the induced population distribution w.r.t. a fixed policy (Guo et al., 2019; Elie et al., 2020; Perrin et al., 2020; Xie et al., 2021; Cui & Koeppl, 2021; Laurière et al., 2022b; Anahtarci et al., 2023). However, FPI-type methods face several limitations: 1) The policy learning and population learning in FPI-type methods typically involves distinct iterative processes, and are implemented separately and executed sequentially, potentially increasing the overall computational burden (Mao et al., 2022; Zaman et al., 2023). 2) Vanilla FPI methods suffer from instability. As the policy or population is fixed while updating the other, the differences between updates in consecutive iterations can be drastic, causing oscillations in the learning process, a phenomenon commonly observed in practice (Cui & Koeppl, 2021). 3) Separating the forward and backward processes prevents us from applying abundant methods developed for policy learning, such as function approximation (Mao et al., 2022; Huang et al., 2024), to population learning.
@@
This work delves into the rapidly growing field of online learning for MFGs (Mao et al., 2022; Angiuli et al., 2022; Zaman et al., 2023; Yardim et al., 2023; Zhang et al., 2024a; Zeng et al., 2025), where an online agent interacts with the environment and gathers observations on a Markov chain to update its policy and population estimate. A key aspect of our approach is recognizing that the Markov chain is jointly parameterized by both the policy and population distribution, which together define the transition kernel. From this perspective, finding the mean field equilibrium (MFE) becomes a problem of identifying the optimal parameter—the fixed point of the Bellman and transition operators. This viewpoint unlocks a large toolbox of stochastic parameter approximation methods for learning MFGs, including simple stochastic gradient descent (SGD)-type methods. Specifically, by treating the policy and population as a unified parameter, we update both estimates fully asynchronously using the same batch of samples, thereby eliminating the need for a forward-backward process.
Building on this perspective, we introduce another powerful technique: linear function approximation (LFA), a widely used approach in stochastic approximation methods for policy learning on large or continuous state-action spaces (Melo et al., 2008; Jin et al., 2020). By applying LFA to the unified parameter—encompassing both the policy and population—we develop the first population-aware LFA (PA-LFA) for learning MFGs on continuous state-action spaces. This is particularly significant, as many common MFGs, such as autonomous driving (Huang et al., 2021; Chen et al., 2023a; b; Mo et al., 2024), flocking (Perrin et al., 2021), and crowd modeling (Lachapelle & Wolfram, 2011; Burger et al., 2013), inherently involve continuous state-action spaces. With PA-LFA, our method enables a simple, fully online, model-free SGD-type method for learning MFGs in these settings, supported by strong theoretical guarantees.
While elegant, treating the policy and population as a unified parameter raises several questions: 1) Updating policy and population estimates fully asynchronously creates a strong coupling between the two, which is absent in forward-backward or multi-loop structures (Yardim et al., 2023; Zhang et al., 2024a) or is alleviated in multi-timescale approaches (Mao et al., 2022; Angiuli et al., 2023; Zaman et al., 2023). How can we analyze this coupling and the non-stationary Markov chain it generates to establish convergence guarantees? 2) Empirically, our method—a simple SGD-type method without additional stabilization mechanisms—outperforms vanilla FPI in stability and efficiency (Section 7). What theoretical explanation underpins this? 3) How can we design PA-LFA for continuous state-action spaces such that each population update maintains comparable operational complexity to that in the finite state-action case? 4) How does PA-LFA influence convergence? Under what conditions does it converge to an MFE, and if it doesn’t, how large is the approximation error?
@@
Main results. With the above questions in mind, we highlight the key contributions of this work:
@@
  • We propose SemiSGD, a simple online SGD-type method for learning MFGs. We innovatively treat the policy and the population as a unified parameter. Algorithmically, we update both simultaneously and asynchronously using the same online observations and learning rate, thus eliminating the forward-backward process typical of FPI methods. Theoretically, we show that the unified parameter in SemiSGD follows a descent direction towards the MFE, whereas neither the policy nor population alone is guaranteed to do so (Lemma 1). This gives a potential explanation for the stability and efficiency of SemiSGD over vanilla FPI (Section 7).
@@
  • Finite-time convergence analysis is provided for SemiSGD with PA-LFA. We novelly regard the learning process as a stochastic approximation on a non-stationary Markov chain parameterized by the unified parameter. This perspective enables a straightforward SGD-type analysis, elegantly handling the strong coupling between the policy and population and offering insights into the learning dynamics of SemiSGD. We prove that, under a contractivity condition no stronger than prior work, SemiSGD converges to the MFE. The contractivity condition can be hard to verify in practice and potentially implies large regularization. In response, we propose a new condition that is more practical. This condition allows general non-regularized policies, under which SemiSGD converges to a neighborhood centered at the MFE. In both scenarios, SemiSGD enjoys state-of-the-art sample complexities.
  • For non-linear MFGs, SemiSGD with PA-LFA converges to the projected MFE (or its neighborhood). We characterize the distance between this projected MFE and the actual MFE in terms of the intrinsic approximation error in the linear representation.
  • We conduct six experiments on three MFGs to demonstrate the properties of our methods: 1) For continuous state-action spaces, PA-LFA is more efficient and accurate than discretization; 2) SemiSGD automatically stabilizes without additional mechanics, achieves a higher accuracy, and is faster by eliminating the forward-backward procedure.
Related work. Learning MFGs has garnered increasing attention, with recent comprehensive reviews by Laurière et al. (2022a); Cui et al. (2022). To address the instability issue of FPI, various stabilization mechanisms have been proposed, broadly categorized into: 1) regularization (Cui & Koeppl, 2021; Anahtarci et al., 2023); 2) fictitious play (Elie et al., 2020; Perrin et al., 2020; Cardaliaguet & Hadikhanloo, 2017); and 3) mirror descent (Perolat et al., 2021; Xie et al., 2021; Yardim et al., 2023; Laurière et al., 2022b). Recently, Angiuli et al. (2022); Zaman et al. (2023); Yardim et al. (2023); Zhang et al. (2024a) developed online oracle-free methods for the population learning in MFGs, where a population estimate is updated using online observations without requiring an oracle to manipulate the population or directly return the population measure. Angiuli et al. (2023); Zaman et al. (2023); Mao et al. (2022); Zeng et al. (2025) further proposed single-loop multi-timescale schemes to asynchronously update the policy and population estimates. On the policy learning side, Mao et al. (2022); Huang et al. (2024) considered function approximation, though they both assume access to the population measure rather than learning it. An extended literature review with detailed comparisons of existing setups and methods is provided in Appendix A.
@@ 2.1 Revisit online learning for MFGs on finite state-action spaces
We consider an infinite-horizon Markov decision process (MDP) denoted by $(\mathcal{S},\mathcal{A},r,P,\gamma)$, with the state space $\mathcal{S}$ and action space $\mathcal{A}$ being finite. In MFGs, the reward function $r$ and transition kernel $P$ depend on the population distribution over the state space. Specifically, for a given state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ and population distribution $\mu\in\mathcal{D}(\mathcal{S})$, $r(s,a,\mu)$ and $P(s^{\prime}\,|\,s,a,\mu)$ denote the reward received and the probability of transitioning to state $s^{\prime}\in\mathcal{S}$. We consider a bounded reward function with $\|r\|_{\infty}\leq R$. $\gamma$ is the discount factor.
An agent in an MFG aims to find a policy $\pi$, which maps a state to a distribution on the action space determining the agent’s actions, that maximizes its expected cumulative discounted reward while interacting with the population. We utilize a value-based approach to calculate policies and assume access to a policy operator $\Gamma_{\pi}$ (Zou et al., 2019; Zhang et al., 2024a) that returns a policy based on an (action-)value function $Q\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. We write $\pi_{Q}\coloneqq\Gamma_{\pi}(Q)$. We define two operators for any value functions $Q,Q^{\prime}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{A}|}$ and population measures $M,M^{\prime}\in\mathcal{D}(\mathcal{S})$:
@@
Then our learning goal, the mean field equilibrium (MFE), is defined as the fixed point of these two operators. With an argmax policy operator, the fixed point of the Bellman operator satisfies the Bellman optimality equation, leading to the standard MFE definition in the reward-maximizing setting (Laurière et al., 2022b; Angiuli et al., 2023); with a regularized policy operator, e.g., softmax, the fixed point corresponds to the regularized MFE (Cui & Koeppl, 2021; Zaman et al., 2023). See Section B.1 for more discussions on our MFE definition.
@@
A typical FPI-type method (approximately) calculates the fixed point of the Bellman operator given the current population distribution, i.e., the best response, and the fixed point of the transition operator given the current value function, i.e., the induced population distribution, alternately. Under certain contractivity conditions, FPI converges to the MFE (Guo et al., 2019).
This work focuses on model-free online learning methods for MFGs. In such methods, an agent maintains estimates about the value function and population measure, and uses online observations to update its estimates. For the value function, temporal difference (TD) methods (Sutton & Barto, 2018) are widely used. Given an online observation tuple $O=(s,a,s^{\prime},a^{\prime})$ and step-size $\alpha$, on-policy TD (or SARSA (Sutton & Barto, 2018)) updates the Q-value function as follows:
@@
For the population distribution, Monte Carlo (MC) sampling (Łatuszyński et al., 2013) is a common choice. Given a new state sample $s^{\prime}$ and step-size $\alpha$, MC updates the population measure as follows:
@@
TD and MC updates are widely used in online learning for MFGs (Angiuli et al., 2022; Zaman et al., 2023; Zhang et al., 2024a). Fixing the policy and population measure (i.e., fixing the Markov kernel), TD converges to the optimal value function, and MC converges to the induced population distribution.
@@ 2.2 Stochastic semi-gradient descent
Combining Equations 2 and 3 in a single pass gives a simple SGD-type method. $\mathfrak{g}$ in Equations 2 and 3 is referred as semi-gradients, as it does not represent the actual gradient of any loss function but provides a similar descent direction to the stationary point, i.e., the zero point of $\mathfrak{g}$ [^1] With a slight abuse of notation, we use a single operator $\mathfrak{g}$ to return semi-gradients throughout the paper, for both the action-value function and population measure; it should be clear from the context and its arguments which one it refers to.We term our method SemiSGD and present it in Algorithm 1, where $(Q_{t};M_{t})$ denotes the concatenated representation of the value function and population measure, as they are both vectors for finite state-action spaces. Similarly, $\mathfrak{g}$ represents the concatenation of the semi-gradients for the value function and population measure.
@@
Algorithm 1 is an online, single-loop, uni-timescale method that is free from any forward-backward process. Not only does it update the value function and population measure estimates in an SGD-like manner, but it provides an actual descent direction through the mean-path semi-gradient, defined as $\bar{\mathfrak{g}}_{(Q,M)}(\cdot)\coloneqq\mathbb{E}_{(Q,M)}\mathfrak{g}(\cdot;O)$, where $O$ is the online observation tuple following the steady distribution induced by the policy and transition kernel determined by $Q$ and $M$.
Lemma 1 (Descent direction; informal).
Suppose $(Q_{*};M_{*})$ is an MFE. Suppose the reward function, transition kernel, and policy operator are Lipschitz continuous with Lipschitz constant $L$. For any value function $Q$ and population measure $M$, with $\Delta Q\coloneqq Q-Q_{*}$ and $\Delta M\coloneqq M-M_{*}$, we have
$\Delta M,\bar{\mathfrak{g}}_{{(Q,M)}}(M)\right>\lesssim$ $-\|\Delta M\|^{2}+L\|\Delta M\|(\|\Delta Q\|+\|\Delta M\|),$ (4)
$\Delta Q,\bar{\mathfrak{g}}_{{(Q,M)}}(Q)\right>\lesssim$ $-\|\Delta Q\|^{2}+L\|\Delta Q\|(\|\Delta Q\|+\|\Delta M\|).$
Due to the coupling between $Q$ and $M$, neither $-\bar{\mathfrak{g}}_{{(Q,M)}}(M)$ nor $-\bar{\mathfrak{g}}_{{(Q,M)}}(Q)$ is guaranteed to be a descent direction, no matter how small the Lipschitz constant $L$ is. However, if $L\lesssim 1/2$, adding the two inequalities gives
${(Q;M)}-{(Q_{*};M_{*})},\bar{\mathfrak{g}}_{{(Q,M)}}({(Q;M)})\right>\lesssim-\|{(Q;M)}-{(Q_{*};M_{*})}\|^{2}/2.$ (5)
@@
A formal and generalized version of Lemma 1 for linear MFGs, which encompasses finite MFGs (Example 2), is proved in Section H.1. We now discuss some key insights from Lemma 1.
[…]
Figure 1: Learning dynamics.
Descent direction. The concatenated semi-gradient points to a descent direction for the concatenated estimate $(Q;M)$, while neither the value function nor population measure alone is guaranteed to follow a descent direction, no matter how small the Lipschitz constant is when far from equilibrium (i.e., large $\|M-M_{*}\|$ and $\|Q-Q_{*}\|$). This strongly suggests treating the value function and population measure as a unified parameter and updating them simultaneously at the same rate. The small Lipschitz constant condition $L\lesssim 1/2$ in the lemma, formalized as Assumption 4 in Section 5, ensures contractivity. Thus, while both methods converge under the contractivity assumption, SemiSGD follows a descent direction but FPI may not, hinting the sample efficiency of SemiSGD. See Figure 1 for an illustration.
Incremental update by design. Incrementally updating parameters or damping update steps, rather than switching to entirely new estimates, is a common stabilization technique in learning MFGs (Laurière et al., 2022a), used in methods like fictitious play and mirror descent (Laurière et al., 2022b). As an SGD-type method, updates in SemiSGD are incremental by design. This, combined with the previous point, suggests that SemiSGD enjoys automatic stabilization without needing additional stabilization mechanisms, as confirmed by our experiments (Section 7). These two insights also offer a potential explanation for the oscillation issues seen in FPI-type methods.
Stochastic approximation on non-stationary Markov chains. The lemma highlights that the value function and population measure jointly shape the landscape of learning MFGs and should be treated as a unified parameter. SemiSGD naturally fits as a stochastic approximation method on a non-stationary Markov chain for finding the equilibrium parameter, with non-stationarity arising from the changing transition kernel as the value function and population measure are updated. Two-timescale approaches are used to mitigate the non-stationarity and parameter coupling (Angiuli et al., 2023; Zaman et al., 2023; Mao et al., 2022). Using a unified parameter, our approach enables a simpler uni-timescale scheme with a straightforward SGD-type analysis that accounts for coupling and achieves better sample complexities (see Table 2). Moreover, for general policy operators, which may fail the small Lipschitz constant condition, the stochastic approximation is on a rapidly changing Markov chain (Zhang et al., 2023). Building on the results for this class of methods, we are able to, for the first time, characterize the finite-time convergence performance of learning MFGs without a contractivity or monotonicity condition.